OpenStudy (anonymous):

The Dimensions of a room are 12 feet by 15 feet by 10 feet. what's the furthest distance between 2 points in the room? PLEASE HELP

OpenStudy (anonymous):

From opposite corner to opposite corner. Pythagorean formula is handy here.

OpenStudy (anonymous):

what''s that formula?

OpenStudy (anonymous):

what shape has the room?

OpenStudy (anonymous):

Rectangular prism room. Pythagorean formula: z^2=x^2+y^2

OpenStudy (anonymous):

not very sure;( don't know how'd do this;(

OpenStudy (anonymous):

12 feet by 15 feet by 10 feet. 3 dimesnsion so it cant be rectangular

OpenStudy (anonymous):

oh! sorry i got it now it is a cube hehe

OpenStudy (anonymous):

ha. alright so next?

OpenStudy (anonymous):

so u must apply the pitagoream theorem for 3 dimensions d=sqrt(x^2 +y^2 +z^2)

OpenStudy (anonymous):

Well, a rectangular prism. The length from one corner to the other: